OpenStudy (yeval76):

Cylinder A has a radius of 12 inches and a height of 6 inches. Cylinder B has a volume of 648π. What is the percent change in volume between cylinders A and B? Cylinder B is 50% smaller than cylinder A. Cylinder B is 25% smaller than cylinder A. Cylinder B is 150% bigger than cylinder A. Cylinder B is 200% bigger than cylinder A.

OpenStudy (yeval76):

@ganeshie8

OpenStudy (yeval76):

@ganeshie8

OpenStudy (mathstudent55):

First, find the volume of cylinder A.

OpenStudy (mathstudent55):

Do you know how to find the volume of a cylinder?

OpenStudy (yeval76):

The volume of cylinder A is 2,712.96

OpenStudy (mathstudent55):

\(\Large V_{cylinder} = \pi r^2 h\) where r = radius of the base, and h = height of the cuylinder

OpenStudy (yeval76):

Yea so the volume of cylinder A is 2,712.96

OpenStudy (yeval76):

So what do I do now?

OpenStudy (mathstudent55):

Correct. Since the volume of cylinder B is written in terms of pi, you can also write the volume of cylinder A in terms of pi. (In other words, don't multiply pi in the formula.) \(\Large V_{A} = 864 \pi\) \(\Large V_B = 648 \pi\) Ok so far?

OpenStudy (yeval76):

Yes mam or sir

OpenStudy (yeval76):

Would you divide Va by Vb and then multiply it by 100...

OpenStudy (mathstudent55):

Now we need the second step of the problem. We need to find the percent change. Look in all choices. Every choice states "Cylinder B is ... than cylinder A." This means we are comparing cylinder B to A.

OpenStudy (mathstudent55):

Here is the formula to calculate percent change: \(percent~change = \dfrac{new~amount - old~amount}{old~amount} \times 100\) If the percent change is positive, it is an increase. If the percent change is negative, it is a decrease. In this case, since we are comparing cylinder B to cylinder A, then old amount = 864 pi new amount = 648 pi (pi will cancel out in the calculation.)

OpenStudy (yeval76):

So will it be 25% smaller then cylinder A right?

OpenStudy (mathstudent55):

Correct.

OpenStudy (yeval76):

Yayyyy Thank You!!! @mathstudent55

OpenStudy (mathstudent55):

You're welcome.